OpenStudy (anonymous):

part a. write an equation of the line that is paralell to the link y = 5x - 4 and contains the point (-3, 6). part b. Write the equation of a line that is perpendicular to the line y = 1/2x and contains the point (2,5)

OpenStudy (jim766):

to write and equation of a line you need a slope and a y intercept parallel lines have the same slope, so if your line is parallel to y = 5x -4 , then your slope must also be 5

OpenStudy (anonymous):

Do i need to do any type of formulas so solve?

OpenStudy (jim766):

if your slope is 5, now all you need is y intercept y = mx + b , the b is the y intercept , substitute in the (-3, 6) and the slope and then solve for b 6 = 5(-3) + b

OpenStudy (jim766):

how do you get the b alone?

OpenStudy (jim766):

what is 5(-3) =

OpenStudy (anonymous):

Would it be -15

OpenStudy (jim766):

right 6 = -15 + b , now how do you get rid of the -15?

OpenStudy (anonymous):

would you add the -15 to the 6, then what you have left is -9b..?

OpenStudy (jim766):

6 = -15 + b +15 +15 21 = b

OpenStudy (jim766):

where did the -9b come from?

OpenStudy (anonymous):

I added 6 to each side... lol

OpenStudy (jim766):

lol ...anyway if your b = 21 and m = 5 all you have to do is put them both into y = mx + b y = 5x + 21

OpenStudy (jim766):

you ready to try b?

OpenStudy (anonymous):

yes sir

OpenStudy (jim766):

lol..

OpenStudy (jim766):

this is the same , but different lol

OpenStudy (jim766):

perpendicular lines do not have the same slopes.... they are opposites (different signs) and reciprocals (flip the fraction if their slope is 1/2 the perpendicular slope will be opposite and reciprocal that one is positive, so mine is negative.... do you know what is the reciprocal of 1/2